OpenStudy (anonymous):

Solve sin(arctan(-1/7))

OpenStudy (anonymous):

-.1414

OpenStudy (anonymous):

And how to do it?

OpenStudy (anonymous):

type it in the calculator

OpenStudy (anonymous):

I need to show the work.

OpenStudy (anonymous):

arctan(-x) = -arc(tanx)

OpenStudy (anonymous):

arc sin(-x) = - arc(sin x)

OpenStudy (anonymous):

sin(-x) = -sinx

OpenStudy (anonymous):

sin(arctan(-1/7))= sin(- arc(tan1/7)) = - sin arc tan/7= - sin arc sin 1/sqrt(50)= -1/sqrt50

OpenStudy (anonymous):

sin(arctan(-1/7))= sin(- arc(tan1/7)) = - sin arc tan1/7= - sin arc sin 1/sqrt(50)= -1/sqrt50

OpenStudy (anonymous):

Where from do you get sqrt50?

OpenStudy (anonymous):

arc tanx = arc sin(1/sqrt(1+x^2))

OpenStudy (anonymous):

Oh, okay. Thanks!

OpenStudy (anonymous):

:)